OpenStudy (anonymous):

y+2/3 = y/x , solve for y pleasee :(

geerky42 (geerky42):

You know how to factor, right?

geerky42 (geerky42):

@xylar7

OpenStudy (anonymous):

(Y+2/3) x= y Xy+ 2/3x = y Xy-y= -2/3x Y(x-1)=-2/3x Y= (-2/3x)/(x-1)